used to raise an object of mass 30 kg. If the plane is 

inclined 5° above the horizontal and the coefficient of friction is 0.20, calculate the ideal 

mechanical advantage, actual mechanical advantage, and the efficiency of this machine.


1
Expert's answer
2021-03-14T19:15:39-0400

Angle of inclination"(\theta)" = "5\degree"

Ideal Mechanical Advantage = "cosec\theta" = 11.47

Since, the coefficient of friction = 0.20

Actual Mechanical Advantage = "\dfrac{cosec\theta}{1+\mu}=\dfrac{11.47}{1.20}=9.56"

Efficiency of the machine = "\dfrac{9.56}{11.47}\times100=83.36\%"